### v1

# ReAcTree: Hierarchical LLM Agent Trees with Control Flow for Long-Horizon Task Planning

Abstract.
Recent advancements in large language models (LLMs) have enabled significant progress in decision-making and task planning for embodied autonomous agents. However, most existing methods still struggle with complex, long-horizon tasks because they rely on a monolithic trajectory that entangles all past decisions and observations, attempting to solve the entire task in a single unified process. To address this limitation, we propose ReAcTree, a hierarchical task-planning method that decomposes a complex goal into more manageable subgoals within a dynamically constructed agent tree. Each subgoal is handled by an LLM agent node capable of reasoning, acting, and further expanding the tree, while control flow nodes coordinate the execution strategies of agent nodes. In addition, we integrate two complementary memory systems: each agent node retrieves goal-specific, subgoal-level examples from episodic memory and shares environment-specific observations through working memory. Experiments on the WAH-NL and ALFRED datasets demonstrate that ReAcTree consistently outperforms strong task-planning baselines such as ReAct across diverse LLMs. Notably, on WAH-NL, ReAcTree achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31%.
Key words and phrases:
Task planning, large language models, decision-making, hierarchical planning, behavior trees, embodied agents
[…]
1. Introduction
Large language models (LLMs) have recently exhibited remarkable reasoning capabilities by generating intermediate reasoning steps Wei et al. (2022); Kojima et al. (2022). These advances open up new possibilities for decision-making and task planning in embodied autonomous agents, with the long-term objective of enabling them to fulfill high-level natural language commands autonomously. As early progress toward this goal, several works Huang et al. (2022a); Ahn et al. (2022); Song et al. (2023); Liang et al. (2023) have demonstrated that LLMs can leverage their pre-trained world knowledge to generate mid-level action sequences without additional training, primarily through prompting rather than handcrafted heuristics or learned policies Eysenbach et al. (2019); Xu et al. (2019).
The decision-making and planning capabilities of LLMs can be further improved by incorporating observations or feedback from external environments. Recent approaches explore adapting the next action Yao et al. (2023); Huang et al. (2022b); Zeng et al. (2023), adjusting the entire plan Shinn et al. (2024), or refining both the plan and the next action Wang et al. (2023a); Sun et al. (2023). However, these methods frequently use a single, monolithic trajectory that entangles all past decisions and observations from multiple subgoals, which increases the risk of hallucination and logical failures Liu et al. (2024b). Another line of research investigates multiple reasoning paths Wang et al. (2023b); Yao et al. (2024); Zhou et al. (2024) for general reasoning tasks. Such methods typically assume the capability to simulate multiple paths and revert to previous states Zhao et al. (2024); Zhuang et al. (2024); Hu et al. (2024), which is rarely feasible in real-world scenarios due to irreversible actions (e.g., slicing an apple) and partial observability.
To tackle these challenges, we introduce ReAcTree, a novel framework that dynamically constructs an agent tree in the subgoal space rather than an action tree over primitive actions. Each LLM-powered agent node is responsible for a subgoal and extends the ReAct paradigm Yao et al. (2023): it can reason, act, or expand the tree by proposing new subgoals with an associated control flow when a task is too complex. Control flow nodes, inspired by Behavior Trees Colledanchise et al. (2018), coordinate these agents by sequencing, fallback, or parallel execution. Conceptually, ReAcTree extends the Least-to-Most prompting strategy Zhou et al. (2023) from static reasoning to dynamic, agentic planning. By decomposing a problem into semantically isolated subgoals and providing each with its own focused context, ReAcTree prevents the propagation of reasoning errors and makes long-horizon tasks more tractable for LLMs. The hierarchical structure of ReAcTree is illustrated in Figure 1.
Figure 1. An illustrative example of how ReAcTree generates an agent tree for the natural language instruction: Please bring one pudding and one juice to the coffee table. The left side shows the hierarchical structure with agent nodes (circles) and control flow nodes (squares); the number inside each circle denotes the execution order, and the attached text box specifies the corresponding subgoal. Control flow nodes are labeled by their types ($\rightarrow$ for sequence, ? for fallback, and $\Rightarrow$ for parallel). The right side highlights the decision-making trajectory of agent node 3, including observation, reasoning, acting, subgoal expansion, and the use of episodic and working memory.
To enhance the in-context learning and coordination capabilities of the proposed agent architecture, we incorporate two complementary memory systems. First, episodic memory assists each agent node in effective in-context learning by providing subgoal-level examples from past experiences that are semantically similar to the current subgoal. It can be bootstrapped from a small set of manually collected trajectories and gradually expanded through successful runs. Second, working memory functions as a shared blackboard, allowing agent nodes to exchange critical observations (e.g., the locations of movable objects). This collective situational awareness reduces redundant searches and mitigates hallucinations. Together, these memory systems enable ReAcTree to learn from the past and coordinate in the present, strengthening the robustness of agentic decision-making under partial observability.
We evaluate ReAcTree by extending LoTa-Bench Choi et al. (2024), a benchmark providing two household task environments (WAH-NL Puig et al. (2021); Choi et al. (2024) with VirtualHome Puig et al. (2018) and ALFRED Shridhar et al. (2020) with AI2THOR Kolve et al. (2017)), to a more challenging partially observable setting. This setup reflects real-world constraints and tests agentic decision-making under uncertainty. Across both environments, ReAcTree consistently outperforms strong baselines across various LLMs. On WAH-NL, it achieves a 61% goal success rate with Qwen 2.5 72B, nearly doubling ReAct’s 31% under the same model size. Notably, even with a much smaller 7B model, it reaches 37%, showing that ReAcTree maintains strong performance despite reduced model capacity. The contributions of its core components are further substantiated by ablation studies on memory systems and control flow types, as well as in-depth analyses of computational cost and failure modes.
In summary, this paper makes the following contributions: (1) ReAcTree, a hierarchical planning framework for long-horizon tasks that dynamically constructs an LLM agent tree, where each agent node independently solves a subgoal while control flow nodes coordinate their execution into a coherent plan; (2) two memory systems: episodic memory that enhances in-context learning ability of agent nodes, and working memory that facilitates information sharing across agent nodes to support collective situational awareness; and (3) extensive experiments on LoTa-Bench under partially observable settings, demonstrating the effectiveness of ReAcTree across strong baselines and providing detailed analysis through ablations, computational cost analysis, and failure case studies. The code is available https://github.com/Choi-JaeWoo/ReAcTree.
Figure 2. Illustration of agent node execution and control flow node execution in ReAcTree.
2. Related Work
LLM-based Embodied Agents. LLMs such as GPT-3, GPT-4, PaLM, LLaMA, and Qwen Brown et al. (2020); Achiam et al. (2023); Chowdhery et al. (2023); Dubey et al. (2024); Yang et al. (2024) have demonstrated impressive few-shot in-context learning capabilities, advancing the state of the art in NLP tasks. Prompting techniques Wei et al. (2022); Kojima et al. (2022); Zhou et al. (2023); Gao et al. (2023) further enhance reasoning by encouraging articulation of intermediate steps. Building on these advancements, research has increasingly explored text-based applications to embodied decision-making. Early studies Huang et al. (2022a); Ahn et al. (2022); Song et al. (2023) showed that LLMs could generate mid-level action sequences without additional training, while later studies introduced code-based plan generation Liang et al. (2023); Singh et al. (2023), the extension of classical agent architectures Ichida et al. (2024), and the integration of environmental feedback or tools Huang et al. (2022b); Zeng et al. (2023); Shen et al. (2024); Lu et al. (2024); Chen et al. (2023). ReAct Yao et al. (2023) enhanced planning by prompting explicit intermediate reasoning, while Reflexion Shinn et al. (2024) applied iterative self-refinement Madaan et al. (2024); Welleck et al. (2023); Paul et al. (2024); Chen et al. (2024b), offering additional flexibility for long-horizon tasks.
Hierarchical Task Planning with LLMs. To tackle more complex, long-horizon tasks, researchers have introduced hierarchical frameworks that decompose goals into manageable planning levels. AdaPlanner Sun et al. (2023) and DEPS Wang et al. (2023a) adopt bi-level hierarchies, refining both an overall plan and next-step decisions with environmental feedback. Other methods combine classical task-and-motion planning with learned low-level controllers guided by LLM reasoning Wong et al. (2023); Liu et al. (2024a); Ding et al. (2023); Shukla et al. (2024); Wang et al. (2025). Some works leverage behavior trees Wang et al. (2024); Chen et al. (2024a); Ao et al. (2025) to organize high-level actions in specialized domains such as human-robot cooking or robotic assembly. However, they typically rely on predefined structures or domain-specific routines. In contrast, our approach, ReAcTree, enables dynamic task decomposition, expanding subgoals in complex environments without being tied to a specific domain.
Tree Search-Based Planning with LLMs. A prominent line of research explores multiple reasoning paths to evaluate diverse hypotheses before committing to a final solution. Such approaches Wang et al. (2023b); Yao et al. (2024); Zhou et al. (2024); Xie et al. (2024); Besta et al. (2024); Hao et al. (2023) show that systematically branching reasoning steps significantly improves performance in reasoning tasks. Several studies extend this idea to agentic planning. For instance, LLM-MCTS Zhao et al. (2024) and ToolChain* Zhuang et al. (2024) search action trees using Monte Carlo Tree Search and A* search, respectively. Tree-Planner Hu et al. (2024) independently samples planning paths and merges them into an action tree, then makes decisions based on grounded observations at each node. However, these methods typically assume reversible actions (i.e., rolling back to prior states) in simulators, which limits their applicability in real-world settings. In contrast, ReAcTree constructs a dynamically expanding LLM agent tree that decomposes complex goals into subgoals, delegates them to agent nodes, and coordinates their execution via behavior tree-inspired control flow, replacing search-based exploration with agent coordination.
3. Preliminaries
Problem Formulation. We consider the agentic planning problem as a sequential decision-making problem aimed at achieving a goal $g$ expressed in natural language. At each time step $t$, the agent has access to the context $c_{t}=(o_{1},a_{1},o_{2},a_{2},\cdots,a_{t-1},o_{t})$, where $o_{i}$ and $a_{i}$ represent the observation and action at time step $i$, respectively. The objective of the agent is to generate the next appropriate action $a_{t}$ based on the context $c_{t}$, with the aim of eventually achieving the goal $g$.
ReAct (Yao et al., 2023). ReAct addresses this problem by interleaving reasoning and action execution using a pre-trained LLM $p_{LLM}$. The action policy is defined as $a_{t}\sim p_{LLM}(\cdot|P,g,c_{t})$, where $P=(P_{sys},P_{ic})$ is the initial prompt, composed of a system prompt $P_{sys}$ and in-context examples $P_{ic}$. The key idea is to use the augmented action space, $\hat{\mathcal{A}}_{t}=\mathcal{A}_{t}\cup\mathcal{L}$, where $\mathcal{A}_{t}$ is the set of executable skills available at time $t$, and $\mathcal{L}$ is the language space representing reasoning steps or thoughts. If $a_{t}\in\mathcal{A}_{t}$, the agent executes the action and obtains a text observation from the environment. If $a_{t}\in\mathcal{L}$, it is called a thought or reasoning trace, which aids in the logical inference of the LLM. In this case, the agent does not receive a new observation from the environment, i.e., $o_{t+1}=\phi$.
4. ReAcTree
ReAcTree is a hierarchical task-planning framework for agentic decision-making in complex environments. It dynamically constructs a tree of agent nodes and control flow nodes, as illustrated in Figure 2. Agent nodes operate as LLM-based task planners that can reason, act, and expand the tree with new subgoals and appropriate control flow. Control flow nodes, inspired by behavior trees (Colledanchise et al., 2018), determine how child agent nodes are executed and how their outcomes are propagated upward in the tree. This design offers two key advantages: (1) isolating subgoals reduces hallucination and logical errors in long trajectories by keeping each agent node focused on its local context and goal, which also enables targeted in-context example selection, and (2) control flow nodes ensure robust and interpretable execution logic for reliable planning.
4.1. ReAcTree Algorithm
ReAcTree algorithm dynamically constructs an agent tree of agent nodes and control flow nodes to accomplish a natural language goal $g$. It begins with a single agent node for the top-level goal and dynamically grows the tree as agents decide to expand. The following describes the execution of agent nodes and control flow nodes.
Agent Nodes. Each agent node $n$ operates as an LLM-based task planner with a specific natural language subgoal $g^{n}$, responsible for decision-making to achieve that goal. At each time step $t$, it accesses its context $c^{n}_{t}=(o^{n}_{1},a^{n}_{1},o^{n}_{2},a^{n}_{2},\dots,a^{n}_{t-1},o^{n}_{t})$, where $o^{n}_{i}$ and $a^{n}_{i}$ represent the observation and action at time step $i$. It then samples an action $a^{n}_{t}$ from an LLM policy $p_{\text{LLM}}(\cdot)$:
$a^{n}_{t}\sim p_{LLM}(\cdot\mid P^{n},g^{n},c^{n}_{t}),$ (1)
where the initial prompt $P^{n}=(P_{sys},P^{n}_{ic})$ consists of a system prompt $P_{sys}$ and agent node-specific in-context examples $P^{n}_{ic}$.
A key feature of ReAcTree is its extended action space, $\hat{\mathcal{A}}^{n}_{t}=\mathcal{A}^{n}_{t}\cup\mathcal{L}\cup\mathcal{E}$, where $\mathcal{A}^{n}_{t}$ represents the set of executable skills at time $t$ (e.g., pick up apple 1); $\mathcal{L}$ is the language space for self-reasoning; and $\mathcal{E}=\mathcal{F}\times\mathcal{L}$ is the expand space, with $\mathcal{F}$ and $\mathcal{L}$ denoting the set of control flow types and the language space used to express subgoals, respectively.
If the action $a^{n}_{t}\in\mathcal{A}^{n}_{t}$ or $a^{n}_{t}\in\mathcal{L}$, the agent operates as in the ReAct framework, either executing actions or engaging in reasoning. If $a^{n}_{t}\in\mathcal{E}$, the agent expands the tree by creating a control flow node as its child and agent nodes for the generated subgoals as grandchildren. Formally, this expansion action is expressed as $a^{n}_{t}=(f^{n},[g^{n}_{1},\dots,g^{n}_{K}])$, where $f^{n}$ is the control flow type and each $g^{n}_{i}$ is a natural language subgoal. A control flow node $n_{f}$ with type $f^{n}$ is then attached as a child of node $n$, and agent nodes $n_{i}$ with subgoals $g^{n}_{i}$ are added as children of $n_{f}$. After expansion, node $n$ executes the control flow node $n_{f}$ and awaits its result. The agent node terminates when one of the following occurs: generating done (returning success), failure, or reaching the maximum decision count (both returning failure).
Control Flow Nodes. Each control flow node coordinates the execution of its child agent nodes according to behavior tree principles (Colledanchise et al., 2018). ReAcTree supports three types of control flow nodes: sequence ($\rightarrow$), fallback (??), and parallel ($\Rightarrow$). The sequence node executes its children sequentially, returning success only if all of them succeed and failing immediately if any child fails. The fallback node also executes children in order but returns success as soon as the first child succeeds and fails only if all children fail. Lastly, the parallel node executes all children sequentially without early termination and aggregates outcomes according to a predefined policy; we adopt majority voting, which is useful for subgoals that can be attempted independently (e.g., placing multiple objects in different locations). After execution, a control flow node reports its overall success or failure to its parent node. The complete pseudocode is provided in Appendix A.
4.2. Memory Systems
To support ReAcTree’s hierarchical planning and decision-making, we introduce two complementary memory systems: episodic memory and working memory. Episodic memory enhances the in-context learning capability of agent nodes by storing and retrieving past subgoal-level experiences. In contrast, working memory facilitates information sharing across nodes by recording environment-specific observations, such as the latest location of movable objects, during task execution.
Episodic Memory. In ReAcTree, each agent node $e$ is responsible for a particular subgoal $g^{e}$, and its entire decision-making trajectory forms a subgoal-level experience. Episodic memory $M_{ep}$ stores experiences from agent nodes involved in task executions that ultimately succeed, even if the individual node did not complete its subgoal.
By storing experiences at the subgoal level, episodic memory maintains shorter, goal-directed trajectories than monolithic planners like ReAct. For instance, while ReAct stores a single long trajectory for Bring pudding and juice to the coffee table, ReAcTree instead stores focused trajectories for each subgoal, such as find and pick up pudding and find and pick up juice. This finer granularity enables more targeted in-context retrieval and improves decision accuracy at each agent node.
Formally, each experience is recorded as a tuple $(t^{e},v^{e},s^{e})$, where: (1) $t^{e}=(g^{e},\,o^{e}_{1},\,a^{e}_{1},\,\dots,\,o^{e}_{T},\,a^{e}_{T})$ denotes the full text trajectory, recording all observations $o^{e}_{t}$ and actions $a^{e}_{t}$ during the node’s lifetime; (2) $v^{e}=f_{\text{sen}}(g^{e})$ is the sentence embedding of the node’s subgoal, computed using a pretrained encoder $f_{\text{sen}}$ such as Sentence-BERT (Reimers and Gurevych, 2019); and (3) $s^{e}\in\{\textit{success},\textit{failure},\textit{expand}\}$ is the termination state of the agent node.
The collected episodic memory $M_{ep}$ provides in-context examples for agent node inference. Before decision-making begins at inference time, an agent node retrieves in-context examples from $M_{ep}$ by comparing its current subgoal $g^{n}$ to the stored goals using cosine similarity. Specifically, the agent embeds its goal $g^{n}$ as $v^{n}=f_{\text{sen}}(g^{n})$ and computes the similarity with each $v^{e}$ in $M_{ep}$:
$\text{sim}(v^{n},\,v^{e})=\frac{v^{n}\cdot v^{e}}{\|\,v^{n}\,\|\;\|\,v^{e}\,\|}.$ (2)
The agent retrieves the top $k$ most similar experiences (with $k$ determined by a token limit) to assemble as in-context examples. If multiple experiences have the same similarity score, ReAcTree samples uniformly across termination states {success, failure, expand} to promote diversity. In practice, episodic memory can be bootstrapped from a few manually collected trajectories and gradually augmented with additional successful executions.
Working Memory. Working memory captures environment-specific information within a single ReAcTree run and is shared across all agent nodes to store and recall key observations. In this work, we focus on tracking the locations of movable objects to reduce redundant interactions and mitigate hallucinations by providing accurate, environment-specific knowledge.
Working memory is integrated into agent nodes of ReAcTree through two mechanisms. First, the executable skill set $\mathcal{A}^{n}_{t}$ is augmented with special actions, recall location of <<movable object>>, enabling the agent to query the stored location of any movable object directly instead of actively exploring the environment. This recall action exclusively queries working memory. Second, working memory is automatically updated whenever an agent observes movable objects during interaction. For instance, if an agent opens a fridge and finds juice, it updates the working memory to reflect that juice is now in the fridge. We implement working memory as a lightweight Python dictionary that maps object classes to lists of observed instances and their associated locations (e.g., IDs, rooms, receptacles). This mechanism extends the concept of tool usage in language models (Schick et al., 2024), where the agent treats recall location as a specialized tool-like action.
Table 1. Performance of ReAcTree, ReAcTree+WM, and 5 baselines across 7 LLMs. Both GSR and SSR (%) are reported. Bold indicates the best result, and underlined indicates the second best. For brevity, we use Phi-4-RP to denote Phi-4-reasoning-plus.
Method LLaMA 3.1 8B LLaMA 3.1 70B Qwen 2.5 7B Qwen 2.5 72B Mistral 7B Gemma 2 9B Phi-4-RP 14B
GSR SSR GSR SSR GSR SSR GSR SSR GSR SSR GSR SSR GSR SSR
ZSP 1.00 13.03 0.00 14.42 0.00 8.98 0.00 14.22 0.00 11.65 1.00 13.87 0.00 17.90
Tree-PlannerN=25 1.00 17.00 2.00 16.72 6.00 22.23 6.00 32.41 1.00 20.43 2.00 17.58 3.00 17.52
Tree-PlannerN=50 4.00 21.85 4.00 23.43 8.00 28.10 9.00 36.03 6.00 23.63 3.00 23.30 4.00 20.40
ReAct 8.00 34.25 30.00 57.05 10.00 31.82 26.00 51.38 6.00 28.18 9.00 37.20 33.00 48.13
ReAct+WM 16.00 42.65 33.00 63.15 13.00 39.73 31.00 54.05 9.00 31.95 11.00 39.93 33.00 51.28
ReAcTree 21.00 51.98 32.00 60.58 18.00 50.20 48.00 75.13 11.00 37.92 26.00 60.43 49.00 67.47
ReAcTree+WM 30.00 60.77 58.00 79.27 37.00 59.63 61.00 79.58 15.00 49.57 38.00 67.08 49.00 69.30
5. Experiments
5.1. Experimental Setup
Datasets and Simulators. We follow the evaluation protocol of LoTa-Bench (Choi et al., 2024), which provides two dataset-simulator pairs for language-based task planning: WAH-NL (Puig et al., 2021; Choi et al., 2024) with VirtualHome (Puig et al., 2018), and ALFRED (Shridhar et al., 2020) with AI2THOR (Kolve et al., 2017). We extend both environments to more realistic, partially observable settings, where the agent receives limited textual observations after each action. All objects additionally include class and instance identifiers for precise grounding. We conduct primary experiments on WAH-NL with VirtualHome, as it features longer-horizon, multi-room tasks with multiple subgoals, and use ALFRED with AI2THOR, which contains relatively short-horizon, single-room, single-goal tasks, mainly for complementary validation.
WAH-NL, an extension of WAH (Puig et al., 2021), comprises 250 training and 100 test tasks across five categories. We use the training set to build episodic memory and the test set for evaluation. To improve reliability, we manually corrected four test tasks with mismatched instructions and goal conditions (see Appendix B.1 for details). ALFRED includes seven task types. In LoTa-Bench, the pick and place two objects task was omitted due to the absence of instance identifiers. We include this task in our evaluation by incorporating simulator-provided instance identifiers. As with WAH-NL, the training set is used to build episodic memory, and evaluation is conducted on the valid-seen and valid-unseen splits.
In both simulators, the agent executes natural language-based primitive actions. After each action, it receives an action-dependent textual observation generated from the simulator’s ground-truth state, revealing only visible objects and receptacles. VirtualHome supports six primitive actions (go to, pick up, put down, open, close, turn on), while AI2THOR supports eight (go to, pick up, put down, open, close, turn on, turn off, slice). Examples of action-observation pairs are provided in Appendix B.2.
Evaluation Metrics. We evaluate task planning performance using two metrics. The goal success rate (GSR) is the percentage of tasks where the agent successfully achieves the overall goal. The subgoal success rate (SSR) is the ratio of completed subgoals to the total number of subgoals. We report both GSR and SSR for WAH-NL, and only GSR for ALFRED, which lacks explicit subgoal definitions.
Implementation Details. We evaluate ReAcTree and its working memory variant, ReAcTree+WM, in a few-shot in-context learning setting without fine-tuning. By default, all experiments include episodic memory unless otherwise specified.
To bootstrap episodic memory, we first manually collected a small number of trajectories and used them as in-context examples to run ReAcTree with LLaMA-3.1 70B (Dubey et al., 2024) over the training set. Only successful executions were retained in episodic memory. For WAH-NL, we collected one manual trajectory per task type and executed the full training set. For ALFRED, we collected three per type and stored up to 100 successful trajectories per type to reduce computational overhead, given the large training set (21K tasks).
The retrieval size was constrained such that the total length of retrieved trajectories used as in-context examples did not exceed 5K tokens. The decision cap was set to 200 for WAH-NL and 100 for ALFRED. We used the Guidance library (Microsoft, 2023) for text generation, employing temperature 0.0 for free-form generation and leveraging its constrained generation capabilities to deterministically select acting actions and control flow types. Prompt templates for both WAH-NL and ALFRED are available in Appendix E.1.
Baselines. We compare ReAcTree and ReAcTree+WM against five baselines: ZSP (Huang et al., 2022a), Tree-Planner (Hu et al., 2024) with $N=25$ or $50$ sampled plans, ReAct (Yao et al., 2023), and its working memory variant ReAct+WM. All baselines are evaluated on WAH-NL using the same retrieval size and decision cap as ReAcTree. For ALFRED, we report results only for ReAct+WM, as it serves as the strongest baseline on WAH-NL. Further implementation details and prompt templates are provided in Appendices C and E.2, respectively.
Node 0 Ensure that both a wine and a juice are on the coffee table. Node 5 Find and pick up the wine in kitchen 1. Node 1 Move the wine onto the coffee table. Node 6 Find and pick up the wine in living room 1. Node 2 Move the juice onto the coffee table. Node 7 Find and pick up the wine in bedroom 1. Node 3 Find and pick up the wine. Node 8 Find and pick up the juice. Node 4 Find the coffee table and put down the wine. Node 9 Find the coffee table and put down the juice.
Figure 3. Success case of ReAcTree on the WAH-NL dataset using Qwen 2.5 72B. The snapshot of each step and the tree structure of nodes are shown. The subgoal of each node is also listed.
5.2. Main Results
We first present the primary evaluation results on the WAH-NL dataset, which features complex, long-horizon tasks under partial observability. Table 1 compares five baselines (ZSP, Tree-Planner with $N=25$ or $50$, ReAct, ReAct+WM) and our methods (ReAcTree, ReAcTree+WM) across seven LLMs ranging from small (7–14B) to large (70–72B), including LLaMA 3.1 (Dubey et al., 2024), Qwen 2.5 (Yang et al., 2024), Mistral (Jiang et al., 2023), Gemma 2 (Team et al., 2024), and Phi-4-reasoning-plus (Abdin et al., 2025) (see the Appendix D for the full list of language models and their HuggingFace identifiers). Both GSR and SSR are reported.
ZSP generates a complete plan at once and thus struggles to adapt under partial observability, resulting in poor performance. Tree-Planner improves this by sampling multiple high-level plans ($N=25$ or $50$) to construct an action tree and adapting its actions based on observations. However, if none of the sampled plans succeed, the system cannot recover. This limitation becomes critical as the search space grows. For example, a typical kitchen may contain multiple kitchen tables, kitchen counters, fridges, and a dishwasher. Without knowing in advance which receptacle contains the target object, plan sampling becomes intractable and frequently fails to generate a valid plan. Even after relaxing action constraints for pick up and put down by omitting the instance identifier (e.g., pick up apple instead of pick up apple 1), the method still yields only marginal improvements over ZSP in our partially observable setting.
Both ReAct and ReAcTree significantly improve GSR and SSR over ZSP and Tree-Planner. While ReAct shows limited gains, ReAcTree consistently achieves higher scores across all LLMs. For example, with Qwen 2.5 72B, ReAcTree+WM achieves a GSR of 61% compared to 31% for ReAct+WM. Its SSR also rises from 54.05% to 79.58%. These consistent gains highlight the effectiveness of decomposing complex tasks into manageable subgoals, where agent nodes handle each semantically isolated subgoal within a focused context and control flow nodes coordinate execution strategies, together making long-horizon tasks more tractable under partial observability.
Notably, ReAcTree enables smaller LLMs to outperform baselines with much larger models. For example, ReAcTree+WM with Qwen 2.5 7B surpasses all baselines in GSR and most in SSR, even against Qwen 2.5 72B and LLaMA 3.1 70B. This improvement stems from two factors: (1) decomposing tasks into simpler subgoals, which keeps the accumulating decision-making trajectory focused on one subgoal at a time, and (2) retrieving subgoal-level examples that are more directly comparable than task-level retrieval. Together, these factors enable smaller models to reason more effectively, substantially narrowing the performance gap across model scales.
Table 2. Performance of ReAct+WM and ReAcTree+WM on the ALFRED across 3 LLMs. GSR (%) is reported for both valid-seen and valid-unseen splits. Bold indicates the best result.
Split Method LLaMA 3.1 Qwen 2.5 Phi-4-RP
8B 70B 7B 72B 14B
Valid- Seen ReAct+WM 21.22 33.31 16.83 37.07 31.71
ReAcTree+WM 25.85 40.00 20.98 40.85 35.12
Valid- Unseen ReAct+WM 19.61 32.40 16.81 39.10 29.72
ReAcTree+WM 26.19 37.03 25.09 39.83 36.18
Table 3. Memory ablation results on Qwen 2.5 7B and 72B. EM, WM describes the memory configuration: none (✗, ✗), WM only (✗, ✓), EM only (✓, ✗), or EM+WM (✓, ✓). We report GSR (%) and SSR (%) with $\Delta$ improvements over the no-memory baseline.
Method EM, WM Qwen 2.5 7B Qwen 2.5 72B
GSR ($\Delta$) SSR ($\Delta$) GSR ($\Delta$) SSR ($\Delta$)
ReAct ✗, ✗  7.00 22.82 13.00 35.75
✗, ✓  6.00 (–1.00) 19.53 (– 3.29) 18.00 (+ 5.00) 44.33 (+ 8.58)
✓, ✗ 10.00 (+3.00) 31.82 (+ 9.00) 26.00 (+13.00) 51.38 (+15.63)
✓, ✓ 13.00 (+6.00) 39.73 (+16.91) 31.00 (+18.00) 54.05 (+18.30)
ReAcTree ✗, ✗  2.00  9.32 31.00 56.82
✗, ✓  1.00 (– 1.00)  7.45 (– 1.87) 47.00 (+16.00) 64.72 (+ 7.90)
✓, ✗ 18.00 (+16.00) 50.20 (+40.88) 48.00 (+17.00) 75.13 (+18.31)
✓, ✓ 37.00 (+35.00) 59.63 (+50.31) 61.00 (+30.00) 79.58 (+22.76)
To demonstrate how ReAcTree plans and executes complex tasks, we analyze a representative task: Make sure there is a wine and a juice on the coffee table. As shown in Figure 3, ReAcTree+WM successfully solves it by decomposing into two subgoals: move the wine onto the coffee table and move the juice onto the coffee table, and executing them in parallel. It explores multiple rooms using fallback strategies, whereas ReAct+WM remains confined to kitchen 1, failing to locate the wine in bedroom 1. Full trajectories are provided in Appendix G, and the corresponding failure case for ReAct+WM is visualized in Appendix F.
We further evaluate ReAcTree+WM on the ALFRED dataset. As shown in Table 2, it consistently outperforms ReAct+WM across all evaluated models and splits. For instance, on the valid-unseen split, it improves over ReAct+WM by 6.58 and 4.63 percentage points with LLaMA 3.1 8B and 70B, respectively, with similar gains observed for Qwen 2.5 and Phi-4-reasoning-plus. These results indicate that ReAcTree remains effective even in relatively short-horizon, single-room tasks and generalizes well to unseen environments.
Beyond quantitative gains, ReAcTree+WM also handles complex tasks that require more precise procedures. For instance, given the instruction place a cooked potato slice in the fridge, both methods slice the potato, but ReAct+WM skips the cooking step and places it directly in the fridge, whereas ReAcTree+WM correctly heats it using the microwave before storage by decomposing the instruction into four subgoals with a sequence control flow node: find and pick up knife, slice and pick up potato, cook and pick up potato, and place potato in fridge. The execution trajectories for both methods are visualized in Appendix H.
5.3. Memory Ablation
All subsequent experiments in this section are conducted on the WAH-NL dataset, featuring long-horizon, multi-room tasks. We study the impact of episodic memory (EM) and working memory (WM) through ablation experiments with Qwen 2.5 models (7B and 72B). When EM is disabled, we use the manually annotated trajectory of a randomly selected training task as a fixed in-context example for all test tasks. For ReAcTree, trajectories from all agent nodes within the selected task are concatenated into a single in-context example. Table 3 reports GSR and SSR of ReAct and ReAcTree under four memory configurations: no memory, WM only, EM only, and EM+WM.
EM and WM show strong synergy. As shown in Table 3, adding either memory component generally improves performance, but combining them consistently yields the highest performance, underscoring their complementary roles. EM provides clean, semantically relevant in-context examples, while WM retains task-relevant observations.
Model scale is a critical factor in EM-disabled settings. The results also highlight a nuanced interaction between memory and model scale, especially when EM is disabled. On the 7B model, both ReAct and ReAcTree perform worse in the WM-only setting than the no-memory baseline. This counterintuitive result occurs because, without EM, the agent receives a fixed and potentially mismatched in-context example. Adding WM further introduces complexity (the recall location of <object> skill) that is not well-grounded by the poor example, leading to confusion for the smaller model with its limited reasoning capacity.
In contrast, the 72B model is robust to this degradation. Its performance improves significantly with WM alone (+16%p GSR for ReAcTree), as its greater reasoning capacity allows it to handle the additional complexity gracefully, even with suboptimal in-context examples.
This dynamic also explains why performance between ReAcTree and ReAct reverses with scale. When EM is disabled, the 7B ReAcTree underperforms because it struggles to parse the complex, multi-granular in-context example, whereas ReAct’s simpler, flat prompt structure is easier for the 7B model to follow. However, at the 72B scale, this trend reverses dramatically: the larger model has sufficient capacity to leverage the hierarchical structure, allowing ReAcTree to outperform ReAct by a large margin (e.g., +18%p in the no-memory setting and +29%p in the WM-only setting).
Hierarchical decomposition is fundamentally powerful. This analysis shows that the hierarchical structure of ReAcTree provides inherent value. The most compelling evidence is that even without memory support, the 72B ReAcTree achieves a GSR of 31.00%, substantially outperforming the 72B ReAct at 13.00%. This confirms that the benefits of hierarchical decomposition are fundamental to the architecture, independent of memory components.
5.4. Control Flow Ablation
We conducted a control flow ablation study of ReAcTree with three configurations: all control flows (all), sequence and fallback (seq+fb), and sequence only (seq). Following the same experimental protocol, we manually collected one trajectory per task type for each setting and bootstrapped additional successful trajectories with LLaMA 3.1 70B to construct episodic memory.
As shown in Table 4, using all consistently yields the best performance, while seq+fb performs comparably. In contrast, relying solely on seq significantly degrades performance. These results highlight the importance of expressive control flows, which coordinate dependencies, enable parallel execution, and recover from subgoal failures in long-horizon tasks.
Table 4. Control flow ablation results. GSR and SSR (%) are reported. Bold indicates the best performance.
Ctrl Flow LLaMA 3.1 8B LLaMA 3.1 70B Qwen 2.5 7B Qwen 2.5 72B
GSR SSR GSR SSR GSR SSR GSR SSR
all 30.00 60.77 58.00 79.27 37.00 59.63 61.00 79.58
seq+fb 28.00 58.77 58.00 78.52 38.00 54.83 61.00 79.08
seq 18.00 45.65 36.00 61.17 15.00 32.18 46.00 63.22
5.5. Computational Cost Analysis
To analyze the computational cost of ReAcTree, we conducted a controlled comparison. We measured three representative configurations using LLaMA 3.1: ReAct+WM (70B), ReAcTree+WM (70B), and ReAcTree+WM (8B). All methods were evaluated on the same 19 tasks that all three configurations successfully completed. For fairness in execution time comparison, all experiments were run on identical hardware with two H100 GPUs and one H200 GPU.
Our analysis considers two perspectives: (1) the trade-off between performance and efficiency, and (2) token-based resource usage as a proxy for GPU memory demand.
First, Table 5 compares the execution time, number of decision steps, and overall performance. At the same 70B model size, ReAcTree+WM requires more decision steps and a longer runtime but achieves a significantly higher GSR (+25%p) than ReAct+WM. This runtime difference stems from a greater number of decision steps rather than a higher computational cost per decision, representing a trade-off for increased robustness and success rate. Notably, ReAcTree+WM (8B) attains performance comparable to ReAct+WM (70B) while being significantly faster, highlighting its efficiency in low-resource settings.
Table 5. Average execution time, decision steps, and goal/subgoal success rates (GSR/SSR) for three configurations on shared successful tasks.
Configuration Time (s) Decision Steps GSR/SSR (%)
ReAct+WM (70B) 109.1 60.1 33 / 62.15
ReAcTree+WM (70B) 198.6 75.2 58 / 79.27
ReAcTree+WM (8B) 69.9 78.0 30 / 60.77
To measure computational resource requirements, we analyzed token usage. Specifically, the maximum number of input tokens serves as an indicator of peak GPU memory demand, while the average number of input/output tokens per decision step reflects the per-step computational load. The results are summarized in Table 6. The average token usage per decision is similar across all models, confirming that the tested configurations of ReAcTree and ReAct incur a comparable computational cost per step. However, ReAct+WM (70B) shows the highest peak input token count (8316) and the largest variance, suggesting greater GPU memory overhead and less predictable resource usage. In contrast, ReAcTree maintains well-bounded and stable token usage due to its modular structure, where each agent node processes a localized subgoal.
Table 6. Peak and average token usage per decision across three configurations.
Configuration Max Input Avg. Input Avg. Output
ReAct+WM (70B) 8316 5359.45 (± 904.83) 16.07 (± 1.95)
ReAcTree+WM (70B) 6977 5362.66 (± 109.06) 17.83 (± 1.50)
ReAcTree+WM (8B) 7173 5390.12 (± 125.72) 17.68 (± 1.73)
5.6. Failure Case Analysis
We analyzed 39 failure cases of ReAcTree+WM with Qwen 2.5 72B in WAH-NL, categorizing them into four types: Ambiguous (10), Execution (12), Search (13), and Expand (4), as visualized in Figure 4.
Ambiguous cases arose from vague instructions (e.g., give me 2 drinks), creating uncertainty about object types or locations. Execution failures, largely due to LLM hallucination, included object confusion (Confusion, 8), irrelevant action repetition (Repeat, 2), and overlooked targets (Skipping, 2), reflecting challenges in reasoning consistency and object recognition. Search failures, the most frequent, resulted from exploration limitations under partial observability, including incomplete searches (PlanFail, 6), local search loops (SameRm, 3), and unnecessary revisits (Revisit, 2). Finally, Expand failures involved incorrect subgoal decomposition, with missing (Missing, 2) or incorrect (Wrong, 2) subgoals.
Search failures were dominant, underscoring the need for stronger fallback and exploration strategies. Addressing Ambiguous cases requires clarifying dialogues, while Execution errors highlight the importance of hallucination control. Less frequent Expand failures point to the need for subgoal refinement in complex scenarios.
Figure 4. Categories and subcategories of failure cases for ReAcTree+WM.
6. Conclusion
We introduced ReAcTree, a hierarchical task planner that extends ReAct by dynamically constructing an agent tree in the subgoal space. Each agent node can reason, act, or expand the tree with new subgoals, while control flow nodes coordinate their execution through sequence, fallback, or parallel strategies. This design prevents the error accumulation typical of single-pass trajectories and makes long-horizon tasks more tractable. To strengthen in-context learning and coordination, we introduced episodic memory, which stores subgoal-level experiences for retrieval, and working memory, which shares environment-specific observations across nodes. Experiments on WAH-NL and ALFRED under partial observability show that ReAcTree consistently outperforms strong baselines across diverse LLMs, demonstrating improved reliability, generalization to unseen environments, and even enabling smaller models to rival larger ones.
Despite these benefits, ReAcTree still faces LLM-specific challenges, including hallucinations and limited capability to recognize failures. It also lacks mechanisms to revise incorrectly expanded subgoals and to handle instruction ambiguity. Future work will focus on mitigating hallucination, refining subgoal correction, and enabling clarification dialogues, ultimately advancing the robustness and deployability of ReAcTree in complex, real-world environments.
Appendix A ReAcTree Pseudocode
This section presents the pseudocode for the two core components of ReAcTree: agent node execution (Algorithm 1) and control flow node execution (Algorithm 2). The planning process begins with a single agent node $n_{0}$ assigned to the top-level goal $g$, which is executed through a call to ExecAgentNode($n_{0}$, $D_{init}=0$), where $D_{init}$ denotes the initial decision count.
At each step, an agent node samples an action from the LLM, selecting from reasoning, acting, and expanding. When expanding, the agent node creates a control flow node as its child. This control flow node then instantiates agent nodes for the generated subgoals as its own children and executes them according to its behavior type, reporting success or failure back to its parent. Algorithms 1 and 2 summarize these procedures.
Algorithm 1 ExecAgentNode($n$, $D_{init}$): Execution of an agent node
1:Input: Agent node $n$ with subgoal $g^{n}$, initial decision count $D_{\text{init}}$, maximum decision count $D_{\max}$
[…]
6:while not terminate do
7:  $a^{n}_{t}\sim p_{\text{LLM}}(\cdot\mid P^{n},g^{n},c^{n}_{t})$ $\triangleright$ Sample next action from LLM
8:  $D\leftarrow D+1$
9:  $t\leftarrow t+1$
10:  if $a^{n}_{t}=\textit{done}$ then
11:   return Success, D
12:  else if $a^{n}_{t}=\textit{failure}$ or $D\geq D_{\max}$ then
13:   return Failure, D
14:  else if $a^{n}_{t}\in\mathcal{A}^{n}_{t}$ then $\triangleright$ Acting: execute an action and receive a textual observation
15:   $o^{n}_{t+1}\leftarrow\textsc{Interact}(a^{n}_{t})$
16:   $c^{n}_{t+1}\leftarrow\textsc{Append}(c^{n}_{t},a^{n}_{t},o^{n}_{t+1})$
17:  else if $a^{n}_{t}\in\mathcal{L}$ then $\triangleright$ Reasoning: produce an internal thought step
18:   $o^{n}_{t+1}\leftarrow\varnothing$
19:   $c^{n}_{t+1}\leftarrow\textsc{Append}(c^{n}_{t},a^{n}_{t},o^{n}_{t+1})$
20:  else if $a^{n}_{t}\in\mathcal{E}$ then $\triangleright$ Expanding: generate control flow and subgoals
21:   $(f^{n},[g^{n}_{1},\dots,g^{n}_{K}])\leftarrow\textsc{ParseExpand}(a^{n}_{t})$
22:   $n_{f}\leftarrow\textsc{InitCtrlFlowNode}(f^{n})$
23:   AddChild($n$, $n_{f}$)
24:   for $i=1$ to $K$ do
25:     $n_{i}\leftarrow\textsc{InitAgentNode}(g^{n}_{i})$
26:     AddChild($n_{f}$, $n_{i}$)
27:   end for
28:   return ExecCtrlFlowNode($n_{f}$, D) $\triangleright$ Execute control flow node (Alg. 2)
29:  end if
30:end while
Algorithm 2 ExecCtrlFlowNode($n_{f}$, $D_{init}$): Execution of a control flow node
1:Input: Control flow node $n_{f}$ with type $f^{n}$ and child agent nodes $\{n_{i}\}$, initial decision count $D_{\text{init}}$
[…]
3:if $f^{n}=\textit{sequence}$ then
4:  for each child node $n_{i}$ in order do
5:   $\mathit{status},D\leftarrow\textsc{ExecAgentNode}(n_{i},D)$ $\triangleright$ Execute agent node (Alg. 1)
6:   if $\mathit{status}=\textit{Failure}$ then
7:     return Failure, D
8:   end if
9:  end for
10:  return Success, D
11:else if $f^{n}=\textit{fallback}$ then
12:  for each child node $n_{i}$ in order do
13:   $\mathit{status},D\leftarrow\textsc{ExecAgentNode}(n_{i},D)$ $\triangleright$ Execute agent node (Alg. 1)
14:   if $\mathit{status}=\textit{Success}$ then
15:     return Success, D
16:   end if
17:  end for
18:  return Failure, D
19:else if $f^{n}=\textit{parallel}$ then
20:  $\mathit{statuses}\leftarrow[\,]$
21:  for each child node $n_{i}$ in order do
22:   $\mathit{status}_{i},D\leftarrow\textsc{ExecAgentNode}(n_{i},D)$ $\triangleright$ Execute agent node (Alg. 1)
23:   $\mathit{statuses}.\text{append}(\mathit{status}_{i})$
24:  end for
25:  $\textit{status}\leftarrow\textsc{AggregatePolicy}(\mathit{statuses})$
26:  return status, D
27:end if
Appendix B Implementation Details for Experimental Setup
B.1. WAH-NL Test Set Corrections
We identified four WAH-NL test tasks where the natural language instructions were inconsistent with the goal conditions, such as referencing incorrect object classes or omitting required items. Therefore, to ensure reliable evaluation, we revised these instructions to better align with their intended goals. Table 7 presents the goal conditions, original instructions, and revised versions.
Table 7. Summary of corrections four WAH-NL test tasks were misaligned with their goal conditions.
Task ID Goal Condition Original Instruction Revised Instruction
4 {‘on_juice_coffeetable’:1, ‘on_wine_coffeetable’:1} Make sure there is a wine and a juice on the kitchen table. Make sure there is a wine and a juice on the coffee table.
8 {‘on_wine_coffeetable’:1, ‘on_apple_coffeetable’:1} Can you please put the apple and wine that are on the coffee table into the fridge please. Can you please put the apple and the wine on the coffee table?
31 {‘inside_waterglass_dishwasher’:1, ‘inside_wineglass_dishwasher’:1, ‘inside_cutleryfork_dishwasher’:1, ‘turnOn_dishwasher’:1} Please, put the cutlery fork and the wine glass in the dishwasher and then turn it on. Please put the cutlery fork, the wine glass, and the water glass in the dishwasher and then turn it on.
94 {‘on_plate_kitchentable’:1, ‘on_waterglass_kitchentable’:1, ‘on_wineglass_kitchentable’:1, ‘on_cutleryfork_kitchentable’:1} Please, put 1 wine glass, 1 water glass, and 1 plate on the table Please put 1 wine glass, 1 water glass, 1 plate, and 1 cutlery fork on the table
B.2. Partially Observable Settings
We extended LoTa-Bench [10] by implementing a rule-based observation generator for both VirtualHome and AI2THOR to support the partially observable setting. After each action, the agent receives a textual description of visible objects and receptacles in its current room, excluding items hidden inside closed containers. These observations are derived from the simulator’s ground-truth state and include class and instance identifiers to enable precise object grounding.
VirtualHome. In VirtualHome, we apply the partially observable protocol, which supports six natural language-based primitive actions (go to, pick up, put down, open, close, and turn on). To capture the variability in observations, we define eight action-dependent observation rules, each specifying how textual feedback is generated based on the semantics of the executed action. Table 8 presents representative action-observation pairs illustrating these rules.
AI2THOR. We apply the same protocol to AI2THOR, which supports eight primitive actions: go to, pick up, put down, slice, open, close, turn on, and turn off. Table 9 presents representative action-observation pairs in AI2THOR, illustrating how textual feedback differs depending on the executed action.
Table 8. Representative action-observation pairs in VirtualHome under the partially observable setting.
Action Context Action Text Observation
initialization - You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bathroom (1). Looking quickly around the room, you see bathroom cabinet (1), bathroom counter (1), faucet (1), sink (1), toilet (1), towel rack (1), washing machine (1).
go to room go to kitchen 1 You move to the kitchen (1). Looking quickly around the room, you see chair (1, 2, 3, 4), dishwasher (1), faucet (1), fridge (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 3), kitchen table (1), microwave oven (1), nightstand (1, 2), sink (1), sofa (1), stove (1), toaster (1).
go to object or receptacle go to kitchen table 1 You arrive at the kitchen table (1). You see apple (2), book (1), bowl (2, 3), cereal (1), chair (1, 2, 3, 4), chocolate syrup (1), condiment bottle (1, 2), condiment shaker (1, 2), cupcake (1, 2), cutlery fork (2, 3, 4), cutlery knife (3, 5, 6), juice (1), kitchen table (1), pancake (1), pear (1), plate (2, 3, 4, 5), pudding (2), water glass (1, 2, 4)
pick up object pick up juice 1 You pick up juice. You hold juice (1).
put down object put down apple 1 You put down apple on coffee table.
open receptacle open fridge 1 You open fridge. You see bench (1), cupcake (1), fridge (1, 2), juice (1).
close receptacle close dishwasher 1 You close dishwasher.
turn on receptacle turn on dishwasher 1 You turn on dishwasher.
Table 9. Representative action-observation pairs in AI2THOR under the partially observable setting.
Action Context Action Text Observation
initialization - You are in the house, and you arrive at the bathroom, Looking quickly around the room, you see TowelHolder (1), Toilet (1), ToiletPaperHanger (1), Sink (1), SinkBasin (1), Shelf (1, 2, 3), HandTowelHolder (1), GarbageCan (1), BathtubBasin (1).
go to object or receptacle go to BathtubBasin (1) You arrive at BathtubBasin (1). You see SoapBar (2), Cloth (1), Bathtub (1), BathtubBasin (1).
pick up object pick up SoapBar (2) You pick up SoapBar (2). You see SoapBar (2), Cloth (1), Bathtub (1), BathtubBasin (1). I am holding SoapBar (2) now.
put down object put down SoapBar (2) You put SoapBar (2) on Shelf (1). You see SoapBar (1, 2, 3), Shelf (1), Candle (2).
slice object slice Potato (2) You slice Potato (2). You see Potato (8, 9, 10, 11, 12), Knife (1), Fridge (1). Potato (5, 6, 7, 8, 9, 10, 11, 12) are sliced. I am holding Knife (1) now.
open receptacle open Microwave (1) You open Microwave (1). You see StoveKnob (1, 2, 3, 4), StoveBurner (3), Potato (1), Pan (1, 2), Microwave (1), Bowl (3).
close receptacle close Microwave (1) You close Microwave (1). You see StoveKnob (1, 2, 3, 4), StoveBurner (3), Pan (1, 2), Microwave (1).
turn on receptacle turn on Microwave (1) You turn on Microwave (1). You see Microwave (1), Drawer (5, 6), Cabinet (7, 8, 10).
turn off receptacle turn off Microwave (1) You turn off Microwave (1). You see Microwave (1), Drawer (5, 6), Cabinet (7, 8, 10).
Appendix C Implementation Details of Baselines
All baseline methods are evaluated under the same few-shot in-context learning setup as ReAcTree, without fine-tuning. They also share identical retrieval size and decision cap: 5K tokens and 200 steps for WAH-NL, and 5K tokens and 100 steps for ALFRED. Details of episodic memory construction, generation strategies, and other implementation aspects are described for each baseline below.
ZSP [19]. For WAH-NL, ZSP constructs episodic memory using a fully observable, rule-based planner to extract 183 ground-truth trajectories from the training set. At test time, it retrieves in-context examples most similar to the current task goal using Sentence-BERT [34], following the shared 5K-token retrieval setup. The LLM performs free-form generation with temperature 0.0 to produce an acting action, which is then mapped to an executable skill via Sentence-BERT-based matching. ZSP generates its full plan recursively without access to intermediate observations, relying solely on the initial environment description. To compensate, we provide global environment information including all rooms, objects, and receptacles at the beginning of each task. Since object perception is unavailable at each step, we relax action constraints by omitting instance identifiers for pick up and put down actions, for example using pick up apple instead of pick up apple 1.
Tree-Planner [18]. For WAH-NL, Tree-Planner follows a two-step procedure consisting of plan sampling and grounded deciding. During plan sampling, it uses the same episodic memory as ZSP (183 rule-based trajectories), retrieves up to 5K tokens of in-context examples, and generates $N$ candidate plans ($N=25$ or $50$) with temperature 0.8 and top-$p=0.95$, following the original paper. Similar to ZSP, it also receives global environment information and generates plans without instance-level identifiers for pick up and put down actions, resulting in an action tree. In the grounded deciding phase, the model receives step-wise observations and selects an action from the pre-generated tree. This decision process uses a fixed set of in-context examples and leverages the Guidance library [30] to perform constrained generation.
ReAct [55] & ReAct+WM. Both methods follow the same implementation settings as ReAcTree in terms of episodic memory construction (manual collection followed by LLM-based bootstrapping), retrieval size and decision cap (5K and 200 for WAH-NL; 5K and 100 for ALFRED), and generation settings using the Guidance library [30] with temperature 0.0 for reasoning and constrained action selection.
Appendix D Language Models
Table 10 lists the specific language models used in our experiments. Model names follow their identifiers in the HuggingFace model hub.
Table 10. Language models used in our experiments.
Class Model name Model size
LLaMA 3.1 meta-llama/Llama-3.1-8B 8B
meta-llama/Llama-3.1-70B 70B
Qwen 2.5 Qwen/Qwen2.5-7B 7B
Qwen/Qwen2.5-72B 72B
Mistral mistralai/Mistral-7B-v0.3 7B
Gemma 2 google/gemma-2-9b 9B
Phi-4-reasonin-plus microsoft/Phi-4-reasoning-plus 14B
Appendix E Prompts Templates
E.1. ReAcTree & ReAcTree+WM
We provide the prompt template of ReAcTree+WM for WAH-NL and ALFRED in Listings 1 and 2, respectively. The only difference between ReAcTree+WM and ReAcTree lies in the omission of the “recall location of” action, while otherwise sharing the same structure. The “” symbol denotes the position where in-context examples retrieved from episodic memory are inserted.
Listing 1: Prompt template of ReAcTree+WM for WAH-NL
You are an advanced robot with ability to think, act, and expand behavior tree nodes in decision-making process. You can perform one of the following tasks:
1. Think: Use reasoning to satisfy the current goal condition.
2. Act: Execute a specific action to accomplish the current goal condition. You should use one of actions of this list: [go to, pick up, put down, open, close, turn on, recall location of, done, failure]
3. Expand: Decompose the current goal condition into more detailed subgoals. When expanding, generate appropriate control flow and subgoals. Control flow can be "sequence" (achieve subgoals sequentially; if any subgoal fails, the sequence is interrupted), "fallback" (attempt subgoals in order until one succeeds; if a subgoal is successful, the remaining subgoals are not attempted), or "parallel" (achieve subgoals in parallel; this enables tasks to continue independently, even if one subgoal fails).
Source domain:
Target_domain:
Your task is to: Put one cupcake and one apple on the coffee table
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (1), cabinet (1), chair (5), desk (1), nightstand (3).
Listing 2: Prompt template of ReAcTree+WM for ALFRED
You are an advanced robot with ability to think, act, and expand behavior tree nodes in decision-making process. You can perform one of the following tasks:
1. Think: Use reasoning to satisfy the current goal condition.
2. Act: Execute a specific action to accomplish the current goal condition. You should use one of actions of this list: [go to, pick up, put down, slice, open, close, turn on, turn off, done, failure]
3. Expand: Decompose the current goal condition into more detailed subgoals. When expanding, generate appropriate control flow and subgoals. Control flow can be "sequence" (achieve subgoals sequentially. If any subgoal fails, the sequence is interrupted) or "fallback" (Attempt subgoals in order until one succeeds. If a subgoal is successful, the remaining subgoals are not attempted).
Source domain:
Target domain:
Your task is to: Hold the clock and turn on the lamp.
You are in the house, and you arrive at the bedroom, Looking quickly around the room, you see Shelf (1, 2, 3, 4, 5, 6), Safe (1), LaundryHamper (1), GarbageCan (1), Drawer (1, 2, 3, 4, 5, 6), Desk (1, 2), Bed (1).
E.2. Baselines
We provide the prompt templates used for the baseline methods ZSP, Tree-Planner, and ReAct+WM across both WAH-NL and ALFRED.
ZSP. ZSP uses the prompt template in Listing 3. The “” portion is replaced by in-context examples relevant to the current task goal.
Tree-Planner. During its plan sampling phase, Tree-Planner employs the same template as ZSP (Listing 3). For grounded deciding, it uses a separate template in Listing 4.
ReAct & ReAct + WM. We provide the prompt templates of ReAct+WM for WAH-NL and ALFRED in Listing 5 and 6, respectively. The only difference from ReAct is the inclusion of the action “recall location of” action from the prompt, while otherwise sharing the same structure. The “” portion in the template indicates where in-context examples retrieved from episodic memory are inserted.
Listing 3: Prompt template of ZSP for WAH-NL
You are an advanced robot with ability to generate action plans. You can perform one of the following actions of this list: [go to, pick up, put down, open, close, turn on, done]
Source domain:
Target_domain:
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). In the kitchen 1, there are chair (1, 2, 3, 4), dishwasher (1), faucet (1), fridge (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 3), kitchen table (1), microwave oven (1), nightstand (1, 2), sink (1), sofa (1), stove (1), toaster (1). In the bathroom 1, there are bathroom cabinet (1), bathroom counter (1), faucet (2), sink (2), toilet (1), washing machine (1). In the bedroom 1, there are bed (1), bookshelf (1), cabinet (1), chair (5), desk (1), nightstand (3). In the living room 1, there are bookshelf (2, 3), chair (6), closet (1), coffee table (1), computer (1), desk (2), nightstand (4), sofa (2, 3), tv (1). The objects in the house are apple (1, 2, 3), bananas (1), bar soap (1), book (1, 2), bowl (1, 2, 3, 4, 5, 6, 7, 8, 9), box (1, 2, 3, 4, 5), bucket (1), candle (1), candy bar (1), cell phone (1, 2), cereal (1), chair (1, 2, 3, 4, 5, 6), chips (1, 2), chocolate syrup (1), coffee pot (1), condiment bottle (1, 2, 3, 4), condiment shaker (1, 2, 3, 4), cooking pot (1, 2), crackers (1, 2), creamy buns (1), cupcake (1, 2, 3), cutlery fork (1, 2, 3, 4), cutlery knife (1, 2, 3, 4, 5, 6), cutlets (1), dishwashing liquid (1), face cream (1, 2, 3), folder (1, 2, 3, 4), frying pan (1), hair product (1, 2, 3, 4), hanger (1, 2, 3, 4, 5, 6, 7), juice (1, 2), keyboard (1), lime (1), milk (1), mouse (1), mug (1, 2, 3), notes (1), oven tray (1), pancake (1), pear (1), pile of clothes (1, 2), pillow (1, 2, 3, 4, 5, 6), plate (1, 2, 3, 4, 5, 6, 7), plum (1, 2), pudding (1, 2), radio (1), rug (1, 2, 3, 4), slice of bread (1, 2), toilet paper (1), toothbrush (1), toothpaste (1), wall phone (1), wall picture frame (1, 2, 3, 4, 5, 6, 7, 8), washing sponge (1), water glass (1, 2, 3, 4, 5), wine (1). You are in the middle of a bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (1), cabinet (1), chair (5), desk (1), nightstand (3).
Your task is to: Put one cupcake and one apple on the coffee table
Listing 4: Prompt template of Tree-Planner grounded deciding for WAH-NL
You need to act as a home robot. At each moment, I will provide you with observations of your current environment, as well as the high-level task I want you to do, and previous mid-level sub-tasks that have been executed. Then, you need to select the best sub-task from the options I provide to complete the designated home task based on the observation and your past experience. When one choosed sub-task causes an error in the environment, you will be provided with the error information and the corresponding sub-task, and you need to re-choose a corrective sub-task at the current time step. For example, the actions that have been executed in the environment are:
go to kitchen 1
go to kitchen table 1
The choosed action is: pick up cutlery fork 1
The prompt (error information) would be: The action: "pick up cutlery fork 1" caused an error: Action is not executable, since the agent is not close to cutlery fork 1 when executing "pick up cutlery fork 1" Among the following actions, wich action would you take.
go to dishwasher 1
go to fridge 1
A corrective choice of action would be: go to dishwasher 1
Currently, you are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (1), cabinet (1), chair (5), desk (1), nightstand (3). You are close to cabinet (1), cell phone (2), desk (1), plate (7). You hold nothing in your hands.
Your task is to: Put one cupcake and one apple on the coffee table
Listing 5: Prompt template of ReAct+WM for WAH-NL
You are an advanced robot with ability to think and act. You can perform one of the following tasks:
1. Think: Use reasoning to satisfy the current goal condition.
2. Act: Execute a specific action to accomplish the current goal condition. You should use one of actions of this list: [go to, pick up, put down, open, close, turn on, recall location of, done, failure]
Source domain:
Target_domain:
Your task is to: Put one cupcake and one apple on the coffee table
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (1), cabinet (1), chair (5), desk (1), nightstand (3).
Listing 6: Prompt template of ReAct+WM for ALFRED
You are an advanced robot with ability to think and act. You can perform one of the following tasks:
1. Think: Use reasoning to satisfy the current goal condition.
2. Act: Execute a specific action to accomplish the current goal condition. You should use one of actions of this list: [go to, pick up, put down, slice, open, close, turn on, turn off, done, failure]
Source domain:
Target domain:
Your task is to: Hold the clock and turn on the lamp.
You are in the house, and you arrive at the bedroom, Looking quickly around the room, you see Shelf (1, 2, 3, 4, 5, 6), Safe (1), LaundryHamper (1), GarbageCan (1), Drawer (1, 2, 3, 4, 5, 6), Desk (1, 2), Bed (1).
Appendix F Failure Case of ReAct+WM
Figure 5. Failure case of ReAct+WM on the WAH-NL dataset using Qwen 2.5 72B.
Appendix G ReAcTree+WM Full Trajectory of Success Case
Listing 7: Full Trajectory of ReAcTree+WM
### Agent Node 0
Your task is to: Make sure there is a wine and a juice on the coffee table.
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bathroom (1). Looking quickly around the room, you see bathroom cabinet (1), bathroom counter (1), faucet (1), sink (1), toilet (1), towel rack (1), washing machine (1).
Think: To solve the task, I need to accomplish 2 subgoals independently. Move the wine on the coffee table, and move the juice on the coffee table.
Expand: {’control_flow’: parallel’, conditions’: move the wine on the coffee table, and move the juice on the coffee table’}
### Agent 1
Your primary goal is to: Make sure there is a wine and a juice on the coffee table.
To achieve this, you should perform your sibling tasks in parallel. At this level, your sibling tasks are: move the wine on the coffee table, and and move the juice on the coffee table.
Your task is to: move the wine on the coffee table
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bathroom (1). Looking quickly around the room, you see bathroom cabinet (1), bathroom counter (1), faucet (1), sink (1), toilet (1), towel rack (1), washing machine (1).
Think: To solve the task, I need to accomplish 2 subgoals sequentially. Find and pick up the wine, and find the coffee table and put down the wine.
Expand: {’control_flow’: sequence’, conditions’: find and pick up the wine, find the coffee table and put down the wine’}
### Agent Node 3
Your primary goal is to: move the wine on the coffee table
To achieve this, you should perform your sibling tasks in sequence. At this level, your sibling tasks are: find and pick up the wine, and find the coffee table and put down the wine.
Your task is to: find and pick up the wine
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bathroom (1). Looking quickly around the room, you see bathroom cabinet (1), bathroom counter (1), faucet (1), sink (1), toilet (1), towel rack (1), washing machine (1).
Think: To solve the task, I need to recall the location of the wine.
Act: recall location of wine
You have not seen wine before.
Think: Now, I need to find and pick up the wine in one of the rooms. The wine is more likely to appear in kitchen 1, living room 1, or bedroom 1. I can check each room one by one until I find and pick up the wine.
Expand: {’control_flow’: fallback’, conditions’: find and pick up the wine in kitchen 1, find and pick up the wine in living room 1, find and pick up the wine in bedroom 1’}
### Agent Node 5
Your primary goal is to: find and pick up the wine
To achieve this, you should perform your sibling tasks using a fallback strategy. At this level, your sibling tasks are: find and pick up the wine in kitchen 1, find and pick up the wine in living room 1, and find and pick up the wine in bedroom 1.
Your task is to: find and pick up the wine in kitchen 1
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bathroom (1). Looking quickly around the room, you see bathroom cabinet (1), bathroom counter (1), faucet (1), sink (1), toilet (1), towel rack (1), washing machine (1).
Think: I am in bathroom 1. Now, I need to go to kitchen 1.
Act: go to kitchen 1
You move to the kitchen (1). Looking quickly around the room, you see bench (1, 2), bookshelf (1, 2), chair (1, 2), dishwasher (1), faucet (2), fridge (1, 2), garbage can (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), kitchen table (1), microwave oven (1), sink (2), stove (1), toaster (1).
Think: Now, I need to find the wine. The wine is more likely to appear on kitchen table (1), or in fridge (1, 2), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8). I can check each place one by one, starting from kitchen table 1, until I find the wine.
Act: go to kitchen table 1
You arrive at the kitchen table (1). You see bench (2), book (1), bowl (6), cereal (1), kitchen table (1), milk (1), pudding (1, 2)
Act: go to fridge 1
You arrive at the fridge (1). The fridge (1) is closed. You see bench (1), cupcake (1), fridge (1)
Act: open fridge 1
You open fridge. You see bench (1), cupcake (1), fridge (1)
Act: go to fridge 2
You arrive at the fridge (2). The fridge (2) is closed. You see bench (1), cupcake (1), fridge (1, 2), juice (1)
Act: open fridge 2
You open fridge. You see bench (1), cupcake (1), fridge (1, 2), juice (1)
Act: go to kitchen cabinet 1
You arrive at the kitchen cabinet (1). The kitchen cabinet (1) is closed. You see chicken (1), cooking pot (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), frying pan (1), garbage can (1), kitchen cabinet (1, 4), kitchen counter (1), sink (2), stove (1)
Act: open kitchen cabinet 1
You open kitchen cabinet. You see chicken (1), cooking pot (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), frying pan (1), garbage can (1), kitchen cabinet (1, 4), kitchen counter (1), sink (2), stove (1)
Act: go to kitchen cabinet 2
You arrive at the kitchen cabinet (2). The kitchen cabinet (2) is closed. You see bowl (1, 2, 3, 4, 5), chair (2), chicken (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), kitchen cabinet (1, 2, 3, 4, 5, 6), kitchen counter (1), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: open kitchen cabinet 2
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chair (2), chicken (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), kitchen cabinet (1, 2, 3, 4, 5, 6), kitchen counter (1), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: go to kitchen cabinet 3
You arrive at the kitchen cabinet (3). The kitchen cabinet (3) is closed. You see bowl (1, 2, 3, 4, 5), chair (1, 2), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: open kitchen cabinet 3
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chair (1, 2), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: go to kitchen cabinet 4
You arrive at the kitchen cabinet (4). The kitchen cabinet (4) is closed. You see bowl (1, 2, 3, 4, 5), chicken (1), cooking pot (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), garbage can (1), kitchen cabinet (1, 2, 4), kitchen counter (1), sink (2), washing sponge (1)
Act: open kitchen cabinet 4
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chicken (1), cooking pot (1), cutlery knife (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), garbage can (1), kitchen cabinet (1, 2, 4), kitchen counter (1), sink (2), washing sponge (1)
Act: go to kitchen cabinet 5
You arrive at the kitchen cabinet (5). The kitchen cabinet (5) is closed. You see bowl (1, 2, 3, 4, 5), chair (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: open kitchen cabinet 5
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chair (1, 2), cutlets (1), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: go to kitchen cabinet 6
You arrive at the kitchen cabinet (6). The kitchen cabinet (6) is closed. You see bowl (1, 2, 3, 4, 5), chair (1, 2), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: open kitchen cabinet 6
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chair (1, 2), dishwashing liquid (1), faucet (2), kitchen cabinet (2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), plate (1, 2, 3, 4), sink (2), washing sponge (1)
Act: go to kitchen cabinet 7
You arrive at the kitchen cabinet (7). The kitchen cabinet (7) is closed. You see bowl (1, 2, 3, 4, 5), chair (1, 2), kitchen cabinet (2, 3, 4, 5, 6, 7, 8), kitchen counter (1), light switch (2), plate (1, 2, 3, 4), sink (2)
Act: open kitchen cabinet 7
You open kitchen cabinet. You see bowl (1, 2, 3, 4, 5), chair (1, 2), kitchen cabinet (2, 3, 4, 5, 6, 7, 8), kitchen counter (1), light switch (2), plate (1, 2, 3, 4), pudding (3), sink (2)
Act: go to kitchen cabinet 8
You arrive at the kitchen cabinet (8). The kitchen cabinet (8) is closed. You see bench (2), chair (1, 2), kitchen cabinet (5, 6, 7, 8), light switch (2), pudding (3)
Act: open kitchen cabinet 8
You open kitchen cabinet. You see bench (2), chair (1, 2), kitchen cabinet (5, 6, 7, 8), light switch (2), pudding (3)
Think: I cannot find the wine in kitchen 1.
Act: failure
### Agent Node 6
Your primary goal is to: find and pick up the wine
To achieve this, you should perform your sibling tasks using a fallback strategy. At this level, your sibling tasks are: find and pick up the wine in kitchen 1, find and pick up the wine in living room 1, and find and pick up the wine in bedroom 1.
Your task is to: find and pick up the wine in living room 1
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a kitchen (1). Looking quickly around the room, you see bench (1, 2), bookshelf (1, 2), chair (1, 2), dishwasher (1), faucet (2), fridge (1, 2), garbage can (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), kitchen table (1), microwave oven (1), sink (2), stove (1), toaster (1).
Think: I am in kitchen 1. Now, I need to go to living room 1.
Act: go to living room 1
You move to the living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1).
Think: Now, I need to find the wine. The wine is more likely to appear on coffee table (1), desk (1), or in closet (1), nightstand (1). I can check each place one by one, starting from coffee table 1, until I find the wine.
Act: go to coffee table 1
You arrive at the coffee table (1). You see alcohol (2), bananas (1), bowl (7), cell phone (1, 2), coffee table (1), lime (1), mug (1), plum (1, 2), sofa (1)
Act: go to desk 1
You arrive at the desk (1). You see chair (3), computer (1), desk (1), nightstand (1), water glass (2)
Act: go to closet 1
You arrive at the closet (1). The closet (1) is closed. You see closet (1), light switch (3)
Act: open closet 1
You open closet. You see closet (1), light switch (3)
Act: go to nightstand 1
You arrive at the nightstand (1). The nightstand (1) is closed. You see box (2, 5, 7), chair (3), computer (1), desk (1), nightstand (1), water glass (2)
Act: open nightstand 1
You open nightstand. You see box (2, 5, 7), chair (3), computer (1), desk (1), nightstand (1), water glass (2)
Think: I cannot find the wine in living room 1.
Act: failure
### Agent Node 7
Your primary goal is to: find and pick up the wine
To achieve this, you should perform your sibling tasks using a fallback strategy. At this level, your sibling tasks are: find and pick up the wine in kitchen 1, find and pick up the wine in living room 1, and find and pick up the wine in bedroom 1.
Your task is to: find and pick up the wine in bedroom 1
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1).
Think: I am in living room 1. Now, I need to go to bedroom 1.
Act: go to bedroom 1
You move to the bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (4), cabinet (1), chair (4), closet (2), coffee table (2), computer (2), desk (2, 3), garbage can (3).
Think: Now, I need to find the wine. The wine is more likely to appear on coffee table (2), or in cabinet (1), closet (2). I can check each place one by one, starting from coffee table 2, until I find the wine.
Act: go to coffee table 2
You arrive at the coffee table (2). You see coffee table (2), magazine (1)
Act: go to cabinet 1
You arrive at the cabinet (1). The cabinet (1) is closed. You see cabinet (1), computer (2), desk (2), light switch (4), sundae (1)
Act: open cabinet 1
You open cabinet. You see cabinet (1), computer (2), desk (2), light switch (4), sundae (1), wine (1)
Think: I find wine 1. Now I need to pick it up.
Act: pick up wine 1
You pick up wine. You hold wine (1).
Think: Now I accomplish the goal.
Act: done
### Agent Node 4
Your primary goal is to: move the wine on the coffee table
To achieve this, you should perform your sibling tasks in sequence. At this level, your sibling tasks are: find and pick up the wine, and find the coffee table and put down the wine.
Your task is to: find the coffee table and put down the wine
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a bedroom (1). Looking quickly around the room, you see bed (1), bookshelf (4), cabinet (1), chair (4), closet (2), coffee table (2), computer (2), desk (2, 3), garbage can (3). You hold wine (1).
Think: I hold wine 1. Now, I need to find the coffee table. The coffee table is more likely to appear in living room 1.
Act: go to living room 1
You move to the living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1). You hold wine (1).
Think: I can see coffee table 1.
Act: go to coffee table 1
You arrive at the coffee table (1). You see alcohol (2), bananas (1), bowl (7), cell phone (1, 2), coffee table (1), lime (1), mug (1), plum (1, 2), sofa (1), wine (1) You hold wine (1).
Think: Now, I need to put down wine 1.
Act: put down wine 1
You put down wine on coffee table
Think: Now I accomplish the goal.
Act: done
### Agent Node 2
Your primary goal is to: Make sure there is a wine and a juice on the coffee table.
To achieve this, you should perform your sibling tasks in parallel. At this level, your sibling tasks are: move the wine on the coffee table, and and move the juice on the coffee table.
Your task is to: and move the juice on the coffee table
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1).
Think: To solve the task, I need to accomplish 2 subgoals sequentially. Find and pick up the juice, and find the coffee table and put down the juice.
Expand: {’control_flow’: sequence’, conditions’: find and pick up the juice, find the coffee table and put down the juice’}
### Agent Node 8
Your primary goal is to: and move the juice on the coffee table
To achieve this, you should perform your sibling tasks in sequence. At this level, your sibling tasks are: find and pick up the juice, and find the coffee table and put down the juice.
Your task is to: find and pick up the juice
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1).
Think: To solve the task, I need to recall the location of the juice
Act: recall location of juice
You saw juice 1 near fridge 2 in kitchen 1.
Think: I can find juice 1 near fridge 2 in kitchen 1. I can see fridge 2.
Act: go to kitchen 1
You move to the kitchen (1). Looking quickly around the room, you see bench (1, 2), bookshelf (1, 2), chair (1, 2), dishwasher (1), faucet (2), fridge (1, 2), garbage can (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), kitchen table (1), microwave oven (1), sink (2), stove (1), toaster (1).
Think: I can see fridge 2.
Act: go to fridge 2
You arrive at the fridge (2). The fridge (2) is open. You see bench (1), cupcake (1), fridge (1, 2), juice (1)
Think: I find juice 1. Now I need to pick it up.
Act: pick up juice 1
You pick up juice. You hold juice (1).
Think: Now I accomplish the goal.
Act: done
### Agent Node 9
Your primary goal is to: and move the juice on the coffee table
To achieve this, you should perform your sibling tasks in sequence. At this level, your sibling tasks are: find and pick up the juice, and find the coffee table and put down the juice.
Your task is to: find the coffee table and put down the juice
You are in the house, and there are 4 rooms: bathroom (1), bedroom (1), kitchen (1), living room (1). You are in the middle of a kitchen (1). Looking quickly around the room, you see bench (1, 2), bookshelf (1, 2), chair (1, 2), dishwasher (1), faucet (2), fridge (1, 2), garbage can (1), kitchen cabinet (1, 2, 3, 4, 5, 6, 7, 8), kitchen counter (1, 2), kitchen table (1), microwave oven (1), sink (2), stove (1), toaster (1). You hold juice (1).
Think: I hold juice 1. Now, I need to find the coffee table. The coffee table is more likely to appear in living room 1.
Act: go to living room 1
You move to the living room (1). Looking quickly around the room, you see bookshelf (3), chair (3), closet (1), coffee table (1), computer (1), desk (1), garbage can (2), nightstand (1), sofa (1), tv (1). You hold juice (1).
Think: I can see coffee table 1.
Act: go to coffee table 1
You arrive at the coffee table (1). You see alcohol (2), bananas (1), bowl (7), cell phone (1, 2), coffee table (1), juice (1), lime (1), mug (1), plum (1, 2), sofa (1), wine (1) You hold juice (1).
Think: Now, I need to put down juice 1.
Act: put down juice 1
You put down juice on coffee table
Think: Now I accomplish the goal.
Act: done
Appendix H Qualitative Results for ALFRED
H.1. Success Case of ReAcTree+WM on ALFRED
Node 0 Place a cooked potato slice in the fridge Node 3 cook and pick up potato Node 1 find and pick up knife Node 4 place potato in fridge Node 2 slice and pick up potato
Figure 6. Success case of ReAcTree+WM on ALFRED using LLaMA 3.1 70B.
H.2. Failure Case of ReAct+WM on ALFRED
Figure 7. Failure case of ReAct+WM on ALFRED using LLaMA 3.1 70B.